OpenStudy (anonymous):

Find the product: (x + 5)(x – 5)

mathslover (mathslover):

(x+5)(x-5) is of the form (a+b)(a-b) And \((a+b)(a-b) = a^2-b^2\) So \((x+5)(x-5)= ?\)

mathslover (mathslover):

@allie_bear22 can u do it now?

OpenStudy (anonymous):

i got x^2 + 25

mathslover (mathslover):

check it again allie. check the signs.

mathslover (mathslover):

\((x+5)(x-5) = x^2 - 5^2 \)

OpenStudy (anonymous):

x^2 − 25

mathslover (mathslover):

Great work, so it is \(x^2-25\) .

OpenStudy (anonymous):

thnx